OpenStudy (anonymous):

One of the tables below contains (x,y) values that were generated by a linear function. Determine which table, and then write the equation of the linear function represented by the table Table 1 x 2 5 8 11 14 17 20 y 1 3 7 13 21 31 43 Table 2 x 1 2 3 4 5 6 7 y 10 13 18 21 26 29 34 Table 3 x 2 4 6 8 10 12 14 y 1 6 11 16 21 26 31

OpenStudy (anonymous):

i desperately need help!

OpenStudy (anonymous):

I believe it's table 3 because the y values are constant now I would chose 2 points and find the slope. then use the point-slope formula to find the equation\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }=m\]for slope and\[y-y _{1}=m(x-x _{1})\] for point-slope